# Protecting information in a parametrically driven hybrid quantum system

### v3

@@ IV Protecting information in the hybrid system
Figure 3 shows the fidelity of the information encoded in the initial state and the temporally evolving photon density matrix $\rho_{c}(t)$, as well as its quasi-probability distribution in terms of $W(p,q)$. Since our motive is to observe decoherence, we start with the maximally coherent initial cavity state |$\ket{\psi_{c}(0)}=\frac{1}{\sqrt{2}}\{\ket{1}+\ket{2}\}$, where |$\ket{1}$ and |$\ket{2}$ are the Fock states in the transformed basis. This is because decoherence effect is the strongest when acting on the maximally coherent state as, loss terms leading to de-excitation such as $\kappa$ and $\gamma_{h}$ are more (less) effective on states with higher (lower) excitation, and the average loss of fidelity is close to that observed for an equal superposition or maximally coherent state. Moreover, effects of loss due to dephasing such as $\gamma_{p}$ is most effective on maximally coherent states. In Fig. 3(a), the plot shows the best-fit envelope of the fidelity of the density matrix at $t$ with the initially prepared superposition state that encodes the information [50]. The decrease in $\mathcal{F}(t)$ corresponds to decoherence or loss of information, as the initially encoded state is no longer accurately accessible. This may take place due to inhomogeneous broadening, as well as other losses in the system. However, under increased parametric driving (higher values of $r$), the loss of fidelity is significantly minimized compared to the undriven system, thus highlighting a robust protection effect on the information encoded in the system. This is also evident in the phase-space properties of the information stored in the cavity as shown in Figs. 3(b)-(d), in terms of the Wigner function of $\rho_{c}(t)$. While Fig. 3(b), shows the $W(p,q)$ for the initially encoded state, Figs. 3(c) and (d) correspond to density matrices after one Rabi cycle for the undriven ($r=0$) and the parametrically driven ($r=2$) system, respectively. These plots clearly show that the phase coherence is better preserved for higher values of $r$.
@@
Notably, for $r=0$, the fidelity drops below ${1}/{2}$ after one Rabi oscillation ($t\sim 1$), which is the probability of randomly guessing the stored information. However, under parametric driving, at large $r$ the fidelity dies down at a rate governed by the losses in the system, while inhomogeneous broadening is suppressed. This highlights the strong protection experienced by the information stored in the hybrid quantum system.
Figure 4: Decay rates (a) $\Gamma_{n}$ of average photon number and (b) $\Gamma_{f}$ of fidelity for inhomogeneous width $\delta=30$ MHz (red triangles) and $\delta=60$ MHz (green dots), obtained from numerical calculations for $N=100$ spins. The solid and dashed curves in (a) correspond to the theoretical value of $\zeta(\tilde{\Omega})$. The shaded region in both plots corresponds to the maximum squeezing of $r=1.38$ (orange) achieved in experiment [43] and $r=1.73$ (gray) in [45]. The inset in (a) shows the suppression in inhomogeneous broadening $\xi_{r}$ calculated from Eq. (5), dot and triangle markers indicates the values computed from numerical simulations for $\delta=30$ and $60$ MHz, respectively. Black cross markers highlights the squeezing $r=1.75$, where inhomogeneity is theoretically suppressed by a factor of $10^{-3}$ for $\delta=30$ MHz.
Figure 4 shows the decay rates of average photon number and fidelity of information encoded in the initial state of the cavity. For $\delta=30$ MHz, we see almost 10 fold reduction in both decay rates for $r=1.5$ and this becomes 20 times at $r=2.25$. Similar reduction in decay rates id observed for $\delta=60$ MHz as well. For $\tilde{\Omega}\gg\delta$, the theoretical decay rates match with the ones obtained numerically and cavity protection effect is visible for large squeezing $r$, with suppression in decay due to inhomogeneous broadening by a factor of $10^{-3}$ at $r=1.75$ for inhomogeneous width of $\delta=30$ MHz (see inset of Fig.  4(a). In accordance with the mean-field calculations, irrespective of the width of the spectral distribution, the decay rate $\Gamma_{n}$ approaches $\kappa+\gamma_{h}+2\gamma_{p}$ in the limit of large $r$. However, the squeezing required to approach this limit increases as the width is increased, because countering high inhomogeneity requires stronger coupling with the cavity [15].
V Discussion
@@
Hybrid quantum systems based on spin ensembles interacting with a cavity are ubiquitous in the design of quantum computing and communication architecture. As such, quantum information stored or transferred using these systems needs to be protected from decoherence arising due to inhomogeneity in the system. While there exist several sophisticated techniques to regain coherence, based on engineering of the spin distribution or using expensive cavities, the present results show that by simply using a parametric drive on the hybrid system and encoding information in the transformed states, an effective protection protocol can be achieved. It is to be noted that this protocol is significantly different from other applications of squeezed light such as its use for enhanced precision measurements [51, 52] or encoding information in continuous variable information processing [53, 54, 55].
@@
From an experimental perspective, the use of higher-order interactions in hybrid systems can be naturally adapted in the design of qubits, registers and quantum memories across a wide range of platforms, ranging from implementation of parametric driving in superconducting circuits [56, 57] and optomechanical devices [58] to the coupling of microwave resonators to ensembles based on electron spins [8], nitrogen-vacancy centers [9], and semiconductor qubits [59]. In recent studies, parametric modulation of potential has been used to observe enhanced effective interaction in trapped ions [41], while superradiant phase transition and strong entanglement has been observed in a nuclear magnetic resonance quantum simulator using antisqueezing effects [60].
@@
Acknowledgements.
We thank Sudipto Singha Roy for helpful suggestions and acknowledge the use of SpaceTime, the high-performance computing facility at IIT Bombay, for a part of the quantum simulations. H.S. acknowledges financial support from the Prime Minister’s Research Fellowship (ID: 1302055), Govt. of India. H.S.D. acknowledges funding from SERB-DST, India under Core-Research Grant (CRG/2021/008918) and IRCC, IIT Bombay (RD/0521-IRCCSH0-001).

### v4

@@ 3 Dynamics in the semiclassical regime
$\frac{d\braket{\hat{a}}}{dt}$ $=-(\kappa+i{\Delta}_{c})\braket{\hat{a}}-i\sum_{k}g_{k}\braket{\hat{\sigma}_{k}^{-}}+i\eta\braket{\hat{a}}^{*},$
$\frac{d\braket{\hat{\sigma}_{k}^{-}}}{dt}$ $=-(\gamma_{h}+2\gamma_{p}+i\Delta_{k})\braket{\hat{\sigma}^{-}_{k}}+ig_{k}\braket{\hat{\sigma}^{z}_{k}}\braket{\hat{a}},$
$\frac{d\braket{\hat{\sigma}_{k}^{z}}}{dt}=-2\gamma_{h}(1+\braket{\hat{\sigma}^{z}_{k}})+2ig_{k}(\braket{\hat{\sigma}_{k}^{-}}\braket{\hat{a}}^{*}-\braket{\hat{\sigma}_{k}^{+}}\braket{\hat{a}}),\\ \leavevmode\nobreak$ (4a)
where $\braket{a^{\dagger}}=\braket{a}^{*}$ and $\braket{\hat{\sigma}_{k}^{+}}=\braket{\hat{\sigma}_{k}^{-}}^{*}$. By solving the equations above, we study the dynamics of a hybrid system consisting of an inhomogeneous ensemble of $N=10^{4}$ spins. This ensemble is sufficiently large to use the semiclassical approximation in the stable region [40], and to drive the system in linear regime, we ensure that $\tilde{g}_{k}\ll\tilde{\Delta}_{c}$. Moreover, to characterize only the effect of inhomogeneous broadening in the semiclassical limit, the photon and spin losses are taken to be very small, as compared to the spin-cavity coupling and width of the distribution.
Figure 2: Semiclassical dynamics of an ensemble of spins inside a cavity. The figures show (a) the time evolution of the average photon number $n(t)=|\braket{\hat{a}_{s}}|^{2}$. The effective coupling is $\Omega\approx 6\times 10^{-4}\leavevmode\nobreak\ \Delta_{c}$, and the spin detuning has a Gaussian frequency distribution of width $\delta=0.75\leavevmode\nobreak\ \Omega$ and mean $\bar{\Delta}_{k}=\tilde{\Delta}_{c}$, and squeezing parameter $r=0.0$ (red dashed), $r=1.0$ (blue dot-dashed), and $r=2.0$ (black dotted), and (b) the variation in decay rate $\Gamma_{n}$ with $r$, for width $\delta=0.75\leavevmode\nobreak\ \Omega$ (red circle), $\delta=1.75\leavevmode\nobreak\ \Omega$ (blue cross), and $\delta=2\leavevmode\nobreak\ \Omega$ (black square). The error bars show the standard deviation between the actual value of $n(t)$ and those corresponding to the fitted decay rate $\Gamma_{n}$. All axes are dimensionless.
Figure 2 captures the dynamics of a hybrid system under parametric driving. The semiclassical equations are derived using the Hamiltonian $\mathcal{H}_{0}$, such that the average photon number in the squeezed frame is given by $n(t)=|\braket{\hat{a}_{s}}|^{2}=\cosh(2r)|\braket{\hat{a}}|^{2}-\sinh(2r)\Re(\hat{a}^{2})$, where $r$ is the squeezing parameter. Similar results are obtained if one instead studies the equations of motion in the squeezed frame, by using $\mathcal{H}=\mathcal{H}_{sq}$ in Eq. (3). In Fig. 2(a), the temporal evolution of $n(t)$ is shown for different values of squeezing parameter $r$, and for a fixed width of the spin detuning distribution $\delta=0.75\leavevmode\nobreak\ \Omega$. It is assumed that the hybrid system is initialized such that the cavity has photon number equal to unity i.e., $n(0)=1$.
The figure shows that for $r=0$, the average photon number quickly decoheres with time due to the inhomogeneity in the spin-ensemble (other sources of dissipation are negligible at this point). However, these excitations are sustained for higher values of $r$, which shows that boosting the parametric drive leads to a reduced decay rate and a longer lifetime for the photon. The rate $\Gamma_{n}$ at which the cavity excitation $n(t)$ is decohering can be captured by fitting the peaks of the Rabi oscillations using the relation, $n(t)=n(0)e^{-\Gamma_{n}t}$. So, higher $\Gamma_{n}$ implies faster decoherence of the photonic excitations in the hybrid quantum system. Figure 2(b) shows the variation of the rate of decay $\Gamma_{n}$ as a function of squeezing parameter $r$, for spin-ensembles of different width $\delta$. The plots clearly show that loss of excitation and decoherence in the system is more severe for ensembles with more inhomogeneity (larger width $\delta$). Importantly, parametric drives of increasing intensities, which give us higher values of $r$, are able to reduce the decay rate $\Gamma_{n}$ successfully even for significantly broadened spin-ensembles, thus giving an overall “cavity protection” effect. In fact, the protection is also observed when the inhomogeneity is larger than the effective coupling, i.e., $\delta>\Omega$, albeit in the absence of any other dissipation in the system.
To better understand the cavity protection effect, the net loss or decay in the system needs to be estimated. This includes losses due to inhomogeneous broadening and losses in the system such as photon decay $\kappa$, spin decay $\gamma_{h}$ and dephasing $\gamma_{p}$. In the semiclassical regime, for large $N$ and $\kappa,\gamma_{h},\gamma_{p}\ll\tilde{\Omega}$, the average photon number in the cavity decays with the rate [15],
$\Gamma_{n}(\tilde{\Omega})=\kappa+\gamma_{h}+2\gamma_{p}+\pi\rho(\tilde{\Omega})\tilde{\Omega}^{2},$ (5)
where $\rho(\tilde{\Omega})\sim 2^{-\tilde{\Omega}^{2}/\delta^{2}}/\delta$, is the Gaussian spectral density of atoms or spins in the ensemble, with width $\delta$. The term $\rho(\tilde{\Omega})\tilde{\Omega}^{2}$ governs the decoherence induced by inhomogeneous broadening, where $\tilde{\Omega}=\Omega\cosh{(r)}$ is the effective coupling strength in the squeezed frame. For a fixed width $\delta$, the cavity protection effect is directly related to $\tilde{\Omega}$, which is a function of the squeezing parameter $r$. For $\tilde{\Omega}\gg\delta$, there is a protective energy gap between the superradiant and the subradiant spin-wave modes, which does not allow inhomogeneity to induce decoherence in the system [14]. As a result, by increasing $r$, the net decay rate $\Gamma_{n}$ can be suppressed and eventually the dynamics of the system becomes independent of inhomogeneous broadening. To quantify the suppression in errors due to inhomogeneity, we define the term
$\xi_{r}\equiv\frac{\Gamma_{n}(\tilde{\Omega})}{\Gamma_{n}(\infty)}-1=\frac{\pi\rho(\tilde{\Omega})\tilde{\Omega}^{2}}{\kappa+\gamma_{h}+2\gamma_{p}}.$ (6)
Note that $\xi_{r}$ becomes negligible in the limit $r\to\infty$ and will therefore serve as a good figure of merit in understanding cavity protection effect in inhomogeneous spin-ensembles.
4 Protecting information in the hybrid system
While semiclassical equations highlight the revival of macroscopic quantities such as the average photon number in the cavity, to study protection of properties such as correlations or quantum information encoded in photonic states, the dynamics of the full quantum system is necessary. This is a computationally challenging task due to the large Hilbert space of the spin-ensemble-cavity system. To overcome this, we use a variational renormalization group method to study driven-dissipative dynamics [41], which is a tensor-network method similar to density matrix renormalization group and matrix-product states [48]. The method employs a time-adaptive approach to temporally evolve the density matrix $\rho$ of a spin-cavity system of about hundred spins, using the Lindblad ME formalism. To achieve this, all states $\rho$ and operators $\hat{O}$ are mapped to a higher dimensional space as superkets $|\rho\rangle\rangle$ and superoperators $\hat{\mathcal{O}}$ i.e., $\rho\rightarrow\textbf{vec}(\rho)=|\rho\rangle\rangle$ and $\hat{O}\rho\rightarrow(\hat{O}\otimes\mathbb{I})|\rho\rangle\rangle=\hat{\mathcal{O}}|\rho\rangle\rangle$. This allows the master equation in Eq. (3) to be mapped to a Schrödinger like equation, $d|\rho\rangle\rangle/dt=\tilde{\mathcal{L}}|\rho\rangle\rangle$, in the superoperator space, where
~$\tilde{\mathcal{L}}=-i(\mathcal{H}\otimes\textbf{I}-\textbf{I}\otimes\mathcal{H}^{T})+\kappa\tilde{\mathcal{L}}_{\hat{a}}\\ +\sum_{k}\{\gamma_{h}\tilde{\mathcal{L}}_{\hat{\sigma}_{k}^{-}}+\gamma_{p}\tilde{\mathcal{L}}_{\hat{\sigma}_{k}^{z}}\},$ (7)
and ~$\tilde{\mathcal{L}}_{\hat{x}}=\hat{x}\otimes\hat{x}^{*}-\frac{1}{2}\hat{x}^{\dagger}\hat{x}\otimes\textbf{I}-\frac{1}{2}\textbf{I}\otimes\hat{x}^{T}\hat{x}$. At this point, the variational renormalization has two key parts. First, finding the renormalized representation for the superket $|\rho\rangle\rangle$ in a significantly reduced subspace, and second, evolving the initial state $|\rho\rangle\rangle$ in a time-adaptive manner. As with most tensor-network methods, the renormalized space is obtained by eliminating the null space or those with marginal singular values. The time evolution is then governed using the Lindblad superoperator in Eq. (7), using the transformed Hamiltonian $\mathcal{H}=\mathcal{H}_{sq}$, and evolved adaptively for small intervals $\Delta t$, such that |$\ket{\rho(t+\Delta t)}\rangle=e^{\tilde{\mathcal{L}}\Delta t}\ket{\rho(t)}\rangle$. Now, ~$\tilde{\mathcal{L}}=\sum_{k}\tilde{\mathcal{L}}_{k}$, and the second-order Suzuki-Trotter decomposition helps expand $e^{\tilde{\mathcal{L}}\Delta t}$ into a product of single spin-cavity evolution: $e^{\tilde{\mathcal{L}}_{N}\Delta t/2}e^{\tilde{\mathcal{L}}_{N-1}\Delta t/2}\dots e^{\tilde{\mathcal{L}}_{1}\Delta t}\dots e^{\tilde{\mathcal{L}}_{N-1}\Delta t/2}e^{\tilde{\mathcal{L}}_{N}\Delta t/2}$ (see Refs. [41, 49]). The computation then follows an iterative sweeping protocol as used in other time-adaptive tensor-network methods [50, 51].
An important impact of parametrically driving an open cavity is that its environment is also affected. In other words, squeezing the cavity also transforms the Lindblad operator $\mathcal{L}_{\hat{a}}$, giving rise to additional noise terms [27] in the master equation. For instance,
$\mathcal{L}_{\hat{a}}[\rho]\rightarrow(n_{\text{th}}+1)\mathcal{L}_{\hat{a}_{s}}[\rho]+n_{\text{th}}\mathcal{L}_{\hat{a}_{s}^{\dagger}}[\rho]\\ -m_{\text{2-ph}}\mathcal{L}^{\prime}_{\hat{a}_{s}}[\rho]-m_{\text{2-ph}}^{*}\mathcal{L}^{\prime}_{\hat{a}_{s}^{\dagger}}[\rho],$ (8)
where $\mathcal{L}^{\prime}_{\hat{x}}[\rho]=\hat{x}\rho\hat{x}-\frac{1}{2}\{\hat{x}^{2},\rho\}$. Here, $n_{\text{th}}$ and $m_{\text{2-ph}}$ are functions of squeezing parameter $r$, and correspond to thermal noise and two-photon correlations introduced by squeezing, respectively. However, these additional errors can be eliminated by coupling the cavity mode to a squeezed-vacuum reservoir with the same squeezing parameter $r$, but phase-shifted by $\pi$ relative to the two-photon drive.
Figure 3: Cavity protection effect in a hybrid system with inhomogeneous spins. The plots show the variation of decay rates $\Gamma_{n}$, for an inhomogeneous ensemble with width $\delta=0.75\leavevmode\nobreak\ \Omega$ (green crosses) and $\delta=1.5\leavevmode\nobreak\ \Omega$ (red triangles), obtained from numerical calculations for $N=100$ spins. The solid-green and dashed-red curves correspond to the theoretical value of $\Gamma_{n}(\tilde{\Omega})$ in the semiclassical limit, with the dotted line marking the theoretical minimum decay rate $\Gamma_{n}(\infty)$. The shaded region corresponds to squeezing regimes $r\leq 1.38$ [36] (orange) and $r\leq 1.73$ [52] (gray) that have been achieved experimentally. The inset shows the variation of $\xi_{r}$, where the markers are same as above. The error bars are the same as in Fig. 2. Both axes are dimensionless.
For the full quantum dynamics of the hybrid system using variational renormalization group, an ensemble of $N=100$ spins is considered inside the cavity. For consistency with semiclassical results, the same set of parameters are used, including the coupling strength $\Omega$. However, we now consider other dissipative terms, such as the photon loss rate $\kappa\approx 10^{-4}\leavevmode\nobreak\ \Delta_{c}$, and radiative loss and dephasing for spins equal to $\gamma_{h}=\kappa/8$ and $\gamma_{p}=\kappa/16$, respectively. The loss rates we consider here are comparable to those observed across different experimental platforms such as transmon qubits [53, 37], trapped ions [34, 35, 36] and NV centres [54, 8, 16, 24]. A detailed comparison of the key parameters is shown in Table 1 in Appendix A.
Figure 3 demonstrates cavity protection effect by highlighting the reduction of the decay rate of average photon number $\Gamma_{n}$ as a function of $\tilde{\Omega}$, the effective coupling strength in the squeezed frame. For $\tilde{\Omega}>\delta$, the numerically obtained values of $\Gamma_{n}$ from the full quantum dynamics of $N=100$ spins are close to the theoretical predictions from the semiclassical limit, for both $\delta=0.75\leavevmode\nobreak\ \Omega$ and $\delta=1.5\leavevmode\nobreak\ \Omega$. The results show a 2.5 fold reduction in $\Gamma_{n}$ for width $\delta\approx 0.75\leavevmode\nobreak\ \Omega$, for squeezing parameter $r=1.5$, and by a factor of 6 for $\delta\approx 1.5\leavevmode\nobreak\ \Omega$ and $r=2.25$. The inset of Fig. 3 exhibits the suppression of error by reducing the factor $\xi_{r}$ by a factor of $10^{3}$ and reducing the decay rate close to $\Gamma_{n}(\infty)$ at $r=1.75$ for $\delta\approx 0.75\leavevmode\nobreak\ \Omega$. Similar reduction is also observed for higher $\delta$, albeit at higher squeezing $r$, which is consistent with the protection effect [15]. The “cavity protection” arising from parametric driving deviates from the semiclassical limit in some aspects. Firstly, for small $r$, the dynamics of the system is affected by the term $g_{k}e^{-r}$ in Eq. (2). Moreover, for $\tilde{\Omega}<\delta$, the system can become weakly coupled, which results in increase in $\Gamma_{n}$ for small $\tilde{\Omega}$ [16], as observed for $\delta=1.5\leavevmode\nobreak\ \Omega$ in Fig. 3. Secondly, for high $r$ and $\tilde{\Omega}$, the decay rate $\Gamma_{n}$ can reduce below the semiclassical lower bound $\Gamma_{n}(\infty)$, as observed for $\delta=0.75\leavevmode\nobreak\ \Omega$. This is due to the finite size of the ensemble (or small $N$), which can lead to the creation of collective dark states as observed in experiments using spectral hole burning [24].
[…]
Figure 4: Dynamics of quantum information encoded in a hybrid system consisting of $N=100$ spins inside a cavity. The width of the frequency distribution is $\delta=0.75\leavevmode\nobreak\ \Omega$. The plots in (a) show the evolution of the fidelity $\mathcal{F}$ between the effective cavity state at time $t$ and the initial state $|\psi_{c}(0)\rangle$ for squeezing parameter $r=0.0$ (dotted black), $r=1.0$ (dot-dashed red), and $r=1.75$ (solid blue). The average fidelity (dashed blue) calculated over 540 Haar random initial states is also shown, with the shaded area representing unit standard deviation. The other subfigures show the Wigner function $W(p,q)$ of the initial cavity state $|\psi_{c}(0)\rangle$ in (b), and the state at $t$ corresponding to fidelity (black triangles in (a)) for parameters (c) $r=0.0$ and (d) $r=1.75$.
@@ 4 Protecting information in the hybrid system
Figure 4 shows the fidelity of the information encoded in the initial state and the temporally evolving photon density matrix $\rho_{c}(t)$, as well as its quasi-probability distribution in terms of $W(p,q)$. The cavity at $t=0$ encodes the superposition state |$\ket{\psi_{c}(0)}=\frac{1}{\sqrt{2}}\{\ket{1}+\ket{2}\}$, where |$\ket{1}$ and |$\ket{2}$ are the energy eigenstates in the transformed basis. While loss in excitations given by $\kappa$ and $\gamma_{h}$ are more (less) effective on states with higher (lower) excitation, dephasing ($\gamma_{p}$) acts strongly on states with higher coherence. As such we consider the maximally coherent state with photon number equal to the mean excitation. In Fig. 4(a), the plot shows the best-fit envelope of the fidelity of the density matrix at $t$ with the initially prepared superposition state that encodes the information [55]. The spin-ensemble has an inhomogeneity given by the width of the distribution $\delta=0.75\leavevmode\nobreak\ \Omega$. As such, the decrease in $\mathcal{F}(t)$ corresponds to decoherence or loss of information, as the initially encoded state is no longer accurately accessible. However, under increased parametric driving ($r=1.0$ and $1.75$), the loss of fidelity is significantly minimized compared to the undriven system ($r=0$), thus highlighting a robust protection effect on the information encoded in the system. This is also evident in the phase-space properties of the information stored in the cavity as shown in Figs. 4(b)-(d), in terms of the Wigner function of $\rho_{c}(t)$. While Fig. 4(b), shows the $W(p,q)$ for the initially encoded state, Figs. 4(c) and (d) correspond to density matrices after one Rabi cycle for the undriven ($r=0$) and the parametrically driven ($r=1.75$) system, respectively. These plots clearly show that the phase coherence is better preserved for higher values of $r$.
@@
Figure 4(a), also shows the evolution of the average fidelity, calculated over 540 Haar random initial states, and its temporal behaviour is quite close to that of the maximally coherent initial state. Notably, for $r=0$, the fidelity drops below ${1}/{2}$ after one Rabi oscillation ($t\sim 1$), which is the probability of randomly guessing the stored information. However, under parametric driving, at large $r$ the fidelity dies down at a rate governed by the losses in the system, while inhomogeneous broadening is suppressed. This highlights the strong protection experienced by the information stored in the hybrid quantum system.
Reference Remarks $\kappa/\omega_{c}$ $\gamma_{h}/\omega_{c}$ $\gamma_{p}/\omega_{c}$ $\delta/\omega_{c}$ $\Omega/\omega_{c}$
Our simulations $10^{-4}$ $\approx 10^{-5}$ $\approx 10^{-6}$ $\approx 10^{-3}$ $\approx 10^{-3}$
Minev et. al. [53] Transmon qubit coupled to SC resonator with two modes $\approx 10^{-4}$ $\approx 10^{-5}$ $\approx 10^{-5}$ $\approx 10^{-3}$
A. Imamoğlu [54] spin-ensemble coupled to SC cavity with a buit-in transmon qubit which provides non-linearity $\approx 10^{-4}$ $\leq 10^{-4}$ $\approx 10^{-3}$
Kubo et.al. [8] NV centre ensemble coupled to SC resonator $\approx 10^{-4}$ $\approx 10^{-3}$ $\approx 10^{-2}$
Putz et. al. [16, 24] NV centre ensemble coupled to SC cavity $\approx 10^{-4}$ $\approx 0$ $\approx 10^{-2}$ $\approx 10^{-2}$
Villiers et. al. [37] Transmon qubit coupled to a waveguide oscillator $\approx 10^{-3}$ $\approx 10^{-3}$ $\approx 10^{-3}$ $\approx 10^{-3}$
Burd et. al. [34] Two Mg2+ ions trapped using radio frequency trap $\approx 10^{-6}$ $\approx 10^{-4}$ $\approx 10^{-3}$
Diniz et. al. [15] Cavity protection effect theory $10^{-4}$ $10^{-7}$ $\approx 10^{-3}$ $\approx 10^{-2}$
Table 1: Comparison of parameters taken in various studies. The experimental demonstration of cavity protection effect is shown in Refs. [16, 24], while the setups in Refs. [34, 37] implement parametric driving. References [53, 54, 8] are based on circuit-QED experiments.
5 Discussion
@@
Hybrid quantum systems based on spin-ensembles interacting with a cavity are ubiquitous in the design of quantum computing and communication architecture. As such, quantum information stored or transferred using these systems needs to be protected from decoherence arising due to inhomogeneity in the system. While there exist several sophisticated techniques to regain coherence, based on engineering of the spin distribution or using expensive cavities, the present results show that by simply using a parametric drive on the hybrid system and encoding information in the transformed states, an effective protection protocol can be achieved.
A pertinent example is the experimental demonstration of cavity protection effect in an NV centre spin-ensemble coupled to a superconducting resonator [16], where a 1.4 times increase $\Omega$ is experimentally achieved. Moreover, it was estimated that a 3-fold increase could be achieved by packing the mode volume with NV centres, but only at the cost of introducing finite spin-spin interaction. In our scheme, the additional enhancement can be achieved using a parametric drive with squeezing factor $r\approx 0.81$, which has been experimentally realized in superconducting circuits [37]. Importantly, this allows strong cavity protection effect without giving rise to spin-spin interactions that may adversely affect the coherent exchange of excitations. Similar parametric driving, with squeezing parameter $r\approx 1$, has also been demonstrated in experiments with trapped ions [35]. Further enhancement can be achieved by increased squeezing up to $r\approx 1.8$, which is close to what can be achieved in state-of-the-art optical experiments [52]. In superconducting quantum interference devices (SQUID), squeezing of $r\approx 2$ has been reported [56].
An important aspect of the method is that it is applicable across different frequency regimes and platforms. For example, parametric driving in platforms such as trapped ions can experimentally achieve squeezing as high as $r\approx 1.4$ [36], using a pumping amplitude $\eta$ in the order of 1 MHz. In experiments with transmon qubit [37], squeezing of $r\approx 0.81$ can be experimentally achieved for $\Delta_{c}/2\pi=20$ MHz and $\eta/2\pi=17$ MHz. While parametric amplification in higher frequency regimes are difficult to achieve in nonlinear materials [57], it is more amenable in superconducting circuits by controlling flux passing through a SQUID [56] or, more recently, using highly anharmonic flux qubits called quartons [58].
@@
It is to be noted that our protocol is significantly different from other applications of squeezed light, such as its use for enhanced precision measurements [59, 60] or encoding information in continuous variable information processing [61, 62, 63]. From an experimental perspective, the use of higher-order interactions in hybrid systems can be naturally adapted in the design of qubits, registers and quantum memories across a wide range of platforms, ranging from implementation of parametric driving in superconducting circuits [64, 65] and optomechanical devices [66] to the coupling of microwave resonators to ensembles based on electron spins [9], nitrogen-vacancy centres [8], and semiconductor qubits [67]. In recent studies, parametric modulation of potential has been used to observe enhanced effective interaction in trapped ions [34], while superradiant phase transition and strong entanglement have been observed in a nuclear magnetic resonance quantum simulator using antisqueezing effects [68].
@@
Note.– The datasets and source code employed in both the mean-field and variational renormalization group methods are available online [69].
We thank Sudipto Singha Roy and Rahul Gupta for helpful suggestions and acknowledge the use of SpaceTime, the high-performance computing facility at IIT Bombay, for a part of the quantum simulations. H.S. acknowledges financial support from the Prime Minister’s Research Fellowship (ID: 1302055), Govt. of India. H.S.D. acknowledges funding from SERB-DST, India under Core-Research Grant (CRG/2021/008918) and IRCC, IIT Bombay (RD/0521-IRCCSH0-001).
Appendix A Simulation Parameters
The parameters used in our mean-field and variational renormalization group methods are comparable to those observed across different hybrid systems. Table 1 compares the cavity decay rates ($\kappa$), spin loss rate ($\gamma_{h}$), spin dephasing rate ($\gamma_{p}$), effective coupling strength ($\Omega$), and inhomogeneity ($\delta$) used in our numerical study alongside values reported in experiments and theory for a host of different platforms. The values chosen in our simulations are of similar orders of magnitude as those in these studies. We note that our framework is agnostic to any specific system, and as such, can be used to simulate the dynamics across all the platforms mentioned in the table.